OpenStudy (anonymous):

7x^2 - 28x + 4y^2 - 8y= -2, hyperbola, ellipse, circle, or parabola?

OpenStudy (anonymous):

After completing the squares, the equation becomes \[7 (x - 2)^2 + 4 (y - 1)^2 = 30 \] It is an ellipse of center (2,1) and with semi axes length \[ \sqrt{\frac {30}{7}}\\ \sqrt{\frac {30}{4}} \]